# DS-Agent: Automated Data Science by Empowering Large Language Models with Case-Based Reasoning

### v3

@@ C.1 Comparison with AutoML Techniques
Evaluation Metric RMLSE (normal-$\downarrow$) MAE (normal-$\downarrow$) Accuracy (normal-$\uparrow$) AUROC (normal-$\uparrow$)
@@ D.1 RankReviser
RankReviser
You are a helpful intelligent system that can identify the informativeness of some cases given a data science problem and experiment log.
Task Description: ‘‘‘
{Here is the task description}
Experiment Log: ‘‘‘
{Here is the experiment log}
Here are some solution cases relevant to this research problem, each indicated by number identifier [].
[1] ‘‘‘
{Here is the first case}
[2] ‘‘‘
{Here is the second case}
[3] ‘‘‘
{Here is the third case}
[4] ‘‘‘
{Here is the fourth case}
[5]‘‘‘
{Here is the fifth case}
Rank 5 cases above based on their relevance, informativess and helpfulness to the task description and the experiment log for planning the next experiment step. The cases should be listed in descending order using identifiers. The most relevant, informative and helpful case should be listed first. The output format should be [] > [], e.g., [1] > [2]. Only response the ranking results, do not say any word or explain.
@@ D.2 Planner
Planner
You are a helpful AI expert assistant, responsible for decision making on the experiment plans. You have the following information including, task description, experiment log, python code, and a relevant case so far.
The task description is:
‘‘‘ Task Description:
{Here is the task description}
The current experiment log is:
‘‘‘ Current experiment log:
{Here is the experiment log}
The python code of last step experiment for the current task description is:
‘‘‘python
{Here is the Python script}
Here is a past experience case written by an human expert for a relevant (but not the same) task:
‘‘‘ Case:
{Here is the retrieved case}
Follow these instructions and do not forget them:
- Incrementally introduce new techniques in your plans to solve the task description, since the programmer who follows your decision cannot handle too many instructions at one time.
- Do not include any technique or step in [Decision] that have been implemented as revealed in the python code.
- Focus on decision making of next single step of experiment. Do not include plans in [Decision] that requires mutiple experiment trials.
- Make sure [Decision] includes all the key points for next step experiment.
- Highlight the supporting experiment results and reasoning before drawing any conclusions.
Make sure that the following prohibitions are not violated:
- Never perform any visualization analysis, since you do not have the ability to view the figures.
- Never change the way of the dataset split in any way during the experiment.
- Never introduce any new features, unless you have enough knowledge of the features and their meanings.
- Never tune more than two hyper-parameters in one experiment step, since this will lead to computation costs.
- Never introduce any technique for distributed training. We only have one single GPU card.
Please carefully reason over this relevant case and the provided task description, and then response exactly in the following format:
[Reflection]: What is the progress of the experiment for this task description? What does the current experiment log and python code reveal?
[Reasoning]: How can the current task description benefit from the relevant case?
[Thought]: To solve this task description and iteratively improve the performance, what is the plans for next experiment trial?
[Check]: List all plans in [Thought] and carefully check (1) whether the plan needs multiple experiment trials; (2) has been implemented in the current python code; or (3) violates the listed prohibitions above.
[Decision]: Give a short, precise but detailed instruction summary on the final experiment plan in next single trial.
@@ D.3 Programmer
Programmer
You are a helpful AI-oriented programming expert. Now, we are solving a data science task. Given this python script:
‘‘‘python
 {Here is the Python script}
Now please edit this script according to the following instructions:
‘‘‘instruction
{Here is the experiment plan}
Note that you should provide the **full** code after the edit, making no other changes. Please ensure the completeness of the codes so that it can be run without additional modifications. Your codes will be executed with the support of a NVIDIA GPU card with 24 GB memory.
Please response exactly in the following format:
‘‘‘python
Here is the python code.
@@ D.4 Debugger
Debugger
You are a helpful AI-oriented programming expert. Now, we are solving a data science task. Given this original python script:
‘‘‘python
{Here is the original Python script}
The instruction for modification is:
‘‘‘instruction
{Here is the experiment plan}
This is the current python code:
‘‘‘python
{Here is the Python script that has bugs}
However, there are some bugs in this version. Here is the execution log:
‘‘‘log
{Here is the output result}
Please revise the script to fix these bugs. Note that you should provide the **full** code after the edit, making no other changes. Please ensure the completeness of the codes so that it can be run without additional modifications. Your codes will be executed with the support of a NVIDIA GPU card with 24 GB memory.
Please response exactly in the following format:
‘‘‘reflection
What leads to error or exception on the last modification version? How to fix it?
‘‘‘python
Provide the corrected python code here.
@@ D.5 Logger
Logger
Given instructions (what is expected to do), execution log (the experimental results) and the code difference (what is actually done and this will be nothing if the experiment failed) of last experiment on the task: {Here is the experiment plan} [Execution Log]: ‘‘‘ {Here is the output result} ‘‘‘ [Code Difference]: ‘‘‘ {Here is the code difference between the codes of this step and last step} ‘‘‘ Here is the running log of your experiment: [Running Log]: ‘‘‘ {Here is the experiment log of the last step} ‘‘‘ Summarize and append the progress of the last step to the running log in this format: [Experiment Summary]: According to the instructions and the code difference, summarize what was experimented in the last step objectively. [Experiment Result]: According to the execution log and the running log, summarize if the last step of experiment brings performance improvement objectively. Only report the performance if this is the first experiment result. Do not include any result that is guessed rather than directly confirmed by the observation. Do not include additional information or suggestions.
@@ D.6 Adapter
Adapter
Here are some example cases that solve data science tasks:
[Task]
{Here is the task description of the case}
[train.py] ‘‘‘python
{Here is the original Python script of the case}
[Solution] ‘‘‘python
{Here is the solution code of the case}
Now please solve the following data science task based on the example cases above.
[Task]
{Here is the task description of the current task}
[train.py] ‘‘‘python
{Here is the original Python script of the current task}
Start the python code with "‘‘‘python". Please ensure the completeness of the code so that it can be run without additional modifications.

### v5

@@ C.1 Comparison with AutoML Techniques
Evaluation Metric RMLSE ($\downarrow$) MAE ($\downarrow$) Accuracy ($\uparrow$) AUROC ($\uparrow$)
@@ C.2.2 Case Study in Deployment Stage
C.3 Error Mode Analyses
In this section, we analyze the traces of DS-Agent and present detailed failure mode analyses for both development stage and deployment stage. Besides, we also present some bad cases at https://github.com/guosyjlu/DS-Agent/blob/main/failure-case.md
C.3.1 Error Mode Analyses in Development Stage
For the development stage, DS-Agent with GPT-4 can achieve 100% success rate due to the strong foundational capabilities of GPT-4, while DS-Agent with GPT-3.5 may fail due to these failure modes:
  • Unreasonable plans, where DS-Agent generates unreasonable plans for the task.
  • Overly complex plans, where DS-Agent generates too complicated plans that Programmer and Debugger cannot accomplish.
  • Failure in debugging, where Debugger cannot successfully rectify errors in the script.
Table 7: Error mode statistics of DS-Agent in the development stage.
Unreasonable Plans Overly Complex Plans Failure in Debugging
DS-Agent (GPT-3.5) 5 2 8
DS-Agent (GPT-4) 0 0 0
As shown in Table 7, DS-Agent with GPT-3.5 tends to fail due to its relatively weaker foundational capabilities. This issue can be mitigated by using stronger LLMs or incorporating further finetuning for better alignment.
C.3.2 Error Mode Analyses in Deployment Stage
For the deployment stage, DS-Agent with GPT-4, GPT-3.5, and Mixtral-8x7b-Instruct present the following failure modes:
  • Shape mismatch, where DS-Agent fails to correctly align shapes within the neural networks.
  • Undefined variables, where DS-Agent attempts to use variables that have not been defined within the script.
  • Incorrect function calling, where DS-Agent incorrectly calls functions with parameters that are not passed properly.
  • Missing package import, where DS-Agent attempts to utilize a package that has not been imported.
  • Key error, where DS-Agent attempts to access keys in the dataframe that do not exist.
  • Data type misalignment, where DS-Agent performs calculations on tensors with incompatible data types.
  • Program incompleteness, where DS-Agent is unable to generate complete programs.
Table 8: Error mode statistics of DS-Agent in the deployment stage.
DS-Agent Mixtral-8x7b-Instruct GPT-3.5 GPT-4
Shape Mismatch 36 17 1
Undefined Variables 7 1 0
Incorrect Function Calling 6 3 0
Missing Package Import 43 5 0
Key Error 4 0 0
Data Type Misalignment 1 0 0
Program Incompleteness 27 0 0
As shown in Table 8, DS-Agent may suffer from bugs with only a single trial in the deployment stage. This issue can be solved by using stronger LLMs, incorporating extra debugging process, etc.
C.4 Some Discussions on DS-Agent
Would it be an issue for the case bank to have a considerable fraction of LLM-generated solutions leading to an echo chamber amplifying LLM biases? There indeed exists a potential risk of amplifying LLM biases when LLM agents reuse LLM-generated solutions to solve tasks. However, this echo chamber can be broken from three perspectives: (1) Since DS-Agent only collects solutions with high performance, the likelihood of collecting low-quality or biased solutions is reduced. Thus, this selective process acts as a quality filter, ensuring that only the most effective solutions are collected. (2) DS-Agent benefits from the CBR loop to iteratively explore and revise the model design based on the execution feedback. This process encourages the discovery of novel combinations of ML techniques, which can help to break the cycle of amplifying existing biases and promote a more diverse set of solutions. (3) LLMs tend to use past ML techniques to organize the solutions. However, since there are so many cutting-edged techniques emerging every day, humans may always propose more innovative solutions that achieve stronger empirical results. As such, LLM-generated solutions cannot considerably cover the leaderboard, and thus, human solutions can be collected.
Dependency on external sources. When external resources are unavailable, DS-Agent can handle this setting via two possible ways: (1) Relying on the intrinsic knowledge of LLMs, which is precisely the ablation variant in Section 4.2.2, DS-Agent w/o CBR. However, the performance of this strategy fully relies on the foundational capabilities of LLMs. (2) Human-AI corporation, where human experts provide the experiment plans and DS-Agent is responsible for implementation. However, this strategy may rely on human experts.
@@ D.1 RankReviser
RankReviser You are a helpful intelligent system that can identify the informativeness of some cases given a data science problem and experiment log. Task Description: ‘‘‘ {Here is the task description} ‘‘‘ Experiment Log: ‘‘‘ {Here is the experiment log} ‘‘‘ Here are some solution cases relevant to this research problem, each indicated by number identifier []. [1] ‘‘‘ {Here is the first case} ‘‘‘ [2] ‘‘‘ {Here is the second case} ‘‘‘ [3] ‘‘‘ {Here is the third case} ‘‘‘ [4] ‘‘‘ {Here is the fourth case} ‘‘‘ [5]‘‘‘ {Here is the fifth case} ‘‘‘ Rank 5 cases above based on their relevance, informativess and helpfulness to the task description and the experiment log for planning the next experiment step. The cases should be listed in descending order using identifiers. The most relevant, informative and helpful case should be listed first. The output format should be [] > [], e.g., [1] > [2]. Only response the ranking results, do not say any word or explain.
@@ D.2 Planner
Planner You are a helpful AI expert assistant, responsible for decision making on the experiment plans. You have the following information including, task description, experiment log, python code, and a relevant case so far. The task description is: ‘‘‘ Task Description: {Here is the task description} ‘‘‘ The current experiment log is: ‘‘‘ Current experiment log: {Here is the experiment log} ‘‘‘ The python code of last step experiment for the current task description is: ‘‘‘python {Here is the Python script} ‘‘‘ Here is a past experience case written by an human expert for a relevant (but not the same) task: ‘‘‘ Case: {Here is the retrieved case} ‘‘‘ Follow these instructions and do not forget them: - Incrementally introduce new techniques in your plans to solve the task description, since the programmer who follows your decision cannot handle too many instructions at one time. - Do not include any technique or step in [Decision] that have been implemented as revealed in the python code. - Focus on decision making of next single step of experiment. Do not include plans in [Decision] that requires mutiple experiment trials. - Make sure [Decision] includes all the key points for next step experiment. - Highlight the supporting experiment results and reasoning before drawing any conclusions. Make sure that the following prohibitions are not violated: - Never perform any visualization analysis, since you do not have the ability to view the figures. - Never change the way of the dataset split in any way during the experiment. - Never introduce any new features, unless you have enough knowledge of the features and their meanings. - Never tune more than two hyper-parameters in one experiment step, since this will lead to computation costs. - Never introduce any technique for distributed training. We only have one single GPU card. Please carefully reason over this relevant case and the provided task description, and then response exactly in the following format: [Reflection]: What is the progress of the experiment for this task description? What does the current experiment log and python code reveal? [Reasoning]: How can the current task description benefit from the relevant case? [Thought]: To solve this task description and iteratively improve the performance, what is the plans for next experiment trial? [Check]: List all plans in [Thought] and carefully check (1) whether the plan needs multiple experiment trials; (2) has been implemented in the current python code; or (3) violates the listed prohibitions above. [Decision]: Give a short, precise but detailed instruction summary on the final experiment plan in next single trial.
@@ D.3 Programmer
Programmer You are a helpful AI-oriented programming expert. Now, we are solving a data science task. Given this python script: ‘‘‘python {Here is the Python script} ‘‘‘ Now please edit this script according to the following instructions: ‘‘‘instruction {Here is the experiment plan} ‘‘‘ Note that you should provide the **full** code after the edit, making no other changes. Please ensure the completeness of the codes so that it can be run without additional modifications. Your codes will be executed with the support of a NVIDIA GPU card with 24 GB memory. Please response exactly in the following format: ‘‘‘python Here is the python code. ‘‘‘
@@ D.4 Debugger
Debugger You are a helpful AI-oriented programming expert. Now, we are solving a data science task. Given this original python script: ‘‘‘python {Here is the original Python script} ‘‘‘ The instruction for modification is: ‘‘‘instruction {Here is the experiment plan} ‘‘‘ This is the current python code: ‘‘‘python {Here is the Python script that has bugs} ‘‘‘‘ However, there are some bugs in this version. Here is the execution log: ‘‘‘log {Here is the output result} ‘‘‘ Please revise the script to fix these bugs. Note that you should provide the **full** code after the edit, making no other changes. Please ensure the completeness of the codes so that it can be run without additional modifications. Your codes will be executed with the support of a NVIDIA GPU card with 24 GB memory. Please response exactly in the following format: ‘‘‘reflection What leads to error or exception on the last modification version? How to fix it? ‘‘‘ ‘‘‘python Provide the corrected python code here. ‘‘‘
@@ D.5 Logger
Logger Given instructions (what is expected to do), execution log (the experimental results) and the code difference (what is actually done and this will be nothing if the experiment failed) of last experiment on the task: {Here is the experiment plan} [Execution Log]: ‘‘‘ {Here is the output result} ‘‘‘ [Code Difference]: ‘‘‘ {Here is the code difference between the codes of this step and last step} ‘‘‘ Here is the running log of your experiment: [Running Log]: ‘‘‘ {Here is the experiment log of the last step} ‘‘‘ Summarize and append the progress of the last step to the running log in this format: [Experiment Summary]: According to the instructions and the code difference, summarize what was experimented in the last step objectively. [Experiment Result]: According to the execution log and the running log, summarize if the last step of experiment brings performance improvement objectively. Only report the performance if this is the first experiment result. Do not include any result that is guessed rather than directly confirmed by the observation. Do not include additional information or suggestions.
@@ D.6 Adapter
Adapter Here are some example cases that solve data science tasks: [Task] {Here is the task description of the case} [train.py] ‘‘‘python {Here is the original Python script of the case} ‘‘‘ [Solution] ‘‘‘python {Here is the solution code of the case} ‘‘‘ Now please solve the following data science task based on the example cases above. [Task] {Here is the task description of the current task} [train.py] ‘‘‘python {Here is the original Python script of the current task} ‘‘‘ Start the python code with "‘‘‘python". Please ensure the completeness of the code so that it can be run without additional modifications.